OpenStudy (anonymous):

Differentiate: f(t) = log_2 (sin t) (with respect to t).

OpenStudy (anonymous):

\[\ln -2sint\] right ?

OpenStudy (anonymous):

No its not -2, it is \[\log_2 (\sin(t))\]

OpenStudy (anonymous):

use this rule the derivation of loga(x)=1 / (x ln(a))